OpenStudy (anonymous):

How do I solve this geometry question? Angle ABD and angle BDE are supplementary. Find the measures of both angles. m angle ABD =5x degrees, m angle BDE = (17x-18)degrees? can someone explain it to me?

OpenStudy (gw2011):

Supplementary means 180 degrees. Therefore, if you add angle ABD and angle BDE they will add to 180 degrees. Here is how you can do this: 5x+17x-18=180 This represents the two angles that add to 180 degrees 22x-18=180 22x=180+18 22x=198 x=9 Substitute this number into the angles: angle ABD=5x angle ABD=(5)(9)=45 degrees angle BDE=17x-18 angle BDE=(17)(9)-18 angle BDE=153-18=135 degrees Therefore, angle ABD=45 degrees and angle BDE=135 degrees

OpenStudy (anonymous):

can you help me with this one? angle ABD and angle BDC are complementary. Find the measures of both angles. m angle ABD = (5y+1)degrees, m angle BDC = (3y-7) degrees

OpenStudy (gw2011):

Complementary means 90 degrees. When you add angle ABD and angle BDC, they will add to 90 degrees. angle ABD+angleBDC=90 degrees 5y+1+3y-7=90 8y-6=90 8y=96 y=12 angle ABD=5y+1 angle ABD=(5)(12)+1 angle ABD=60+1 angle ABD=61 degrees angle BDC=3y-7 angle BDC=(3)(12)-7 angle BDC=36-7 angle BDC=29 degrees